# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ F.5.3 Preprocessing time per neighborhood
Graph Tasks Topological Tasks Nbhd MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 2.42 35.15 53.37 53.82 142.01 333.65 333.65 333.65 Adj-2 3.24 49.49 72.94 72.80 195.71 823.68 823.68 823.68 Adj-3 3.26 58.49 74.16 73.40 199.37 973.49 973.49 973.49 Inc-1 3.62 79.37 92.40 92.20 229.92 1121.78 1121.78 1121.78 Inc-2 4.39 80.91 111.65 111.49 280.18 1544.42 1544.42 1544.42 Inc-3 5.58 123.97 147.31 148.10 367.95 2233.27 2233.27 2233.27 Mix-1 7.27 165.71 191.94 193.45 489.55 3724.43 3724.43 3724.43 Mix-2 4.99 95.34 118.46 117.35 318.12 2495.14 2495.14 2495.14 HOPSE-G Adj-1 4.54 45.94 103.91 103.47 297.61 913.38 913.38 913.38 Adj-2 7.81 65.54 172.20 172.97 496.73 1676.79 1676.79 1676.79 Adj-3 7.76 92.15 174.23 174.49 502.85 2742.26 2742.26 2742.26 Inc-1 8.05 93.32 181.82 183.44 532.93 2654.36 2654.36 2654.36 Inc-2 11.21 90.87 245.87 248.52 714.05 2813.64 2813.64 2813.64 Inc-3 14.53 142.22 324.62 326.91 952.95 4432.76 4432.76 4432.76 Mix-1 21.02 210.41 474.58 472.70 1381.15 7329.51 7329.51 7329.51 Mix-2 14.28 160.55 314.86 319.78 932.91 5241.30 5241.30 5241.30
Table 13: Preprocessing time per neighborhood on cell. Results are shown as mean time in seconds. The best result per domain and model is in bold. Similarly to the case of simplicial, the fastest neighborhoods to precompute are the smallest ones.

### v2

@@ I.1 Total runtime
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 $9.62\pm 0.00$.40) $49.15\pm 0.39$.24) $160.18\pm 0.29$.76) $145.63\pm 0.13$.40) $1488.50\pm 2.83$.52) $1754.92\pm 11.74$.29) $2281.97\pm 4.78$.17) $2498.81\pm 1.86$.09) Adj-2 $10.97\pm 0.01$.91) $49.32\pm 0.16$.14) $141.54\pm 0.34$.29) $177.84\pm 0.49$.45) $2855.15\pm 4.79$.89) $2455.75\pm 4.27$.45) $2478.94\pm 5.20$.52) $2331.82\pm 3.46$.01) Adj-3 $12.66\pm 0.02$.91) $47.95\pm 0.23$.11) $225.36\pm 0.45$.43) $195.86\pm 0.36$.31) $599.04\pm 0.42.11$) $3576.27\pm 3.02$.11) $3043.11\pm 8.31$.36) $2561.17\pm 1.63$.73) Inc-1 $11.50\pm 0.02$.44) $58.85\pm 0.13$.74) $169.46\pm 0.42$.22) $220.22\pm 0.32$.95) $2655.44\pm 4.36$.38) $2963.63\pm 5.10$.20) $2287.67\pm 3.48$.83) $2613.46\pm 1.72$.92) Inc-2 $10.65\pm 0.01$.13) $44.15\pm 0.20$.68) $162.13\pm 0.44$.12) $139.93\pm 0.22$.47) $1923.95\pm 3.35$.87) $3287.92\pm 4.24$.70) $2439.12\pm 4.94$.99) $2178.14\pm 3.33$.52) Inc-3 $11.20\pm 0.01$.06) $61.33\pm 0.21$.24) $236.34\pm 0.21$.93) $194.27\pm 0.37$.37) $2620.02\pm 7.80$.01) $3292.73\pm 4.44$.17) $2825.51\pm 4.50$.02) $2386.66\pm 3.17$.80) Mix-1 $12.16\pm 0.00$.93) $74.15\pm 0.14$.39) $372.57\pm 0.43$.27) $235.41\pm 0.44$.74) $3168.39\pm 5.57$.44) $4205.60\pm 15.93$.84) $6681.04\pm 12.61$.16) $3591.07\pm 4.58$.88) Mix-2 $11.17\pm 0.00$.84) $57.38\pm 0.19$.99) $210.32\pm 0.53$.70) $217.65\pm 0.27$.40) $1630.30\pm 0.75$.76) $3497.96\pm 11.59$.55) $3737.92\pm 2.34$.34) $3590.55\pm 4.68$.47) HOPSE-G Adj-1 $4.56\pm 0.00$.62) $25.49\pm 0.06$.16) $92.97\pm 0.27$.48) $104.94\pm 0.28$.21) $542.03\pm 0.94$.68) $1891.07\pm 0.17$.81) $1885.35\pm 0.28$.06) $2194.20\pm 0.42$.70) Adj-2 $4.92\pm 0.00$.37) $32.03\pm 0.09$.91) $118.32\pm 0.17$.84) $111.18\pm 0.18$.27) $480.02\pm 1.00$.70) $2172.54\pm 2.16$.62) $2096.15\pm 0.33$.94) $2323.75\pm 2.22$.16) Adj-3 $5.58\pm 0.01$.28) $39.85\pm 0.17$.90) $120.88\pm 0.31$.47) $134.65\pm 0.49$.81) $478.06\pm 1.07$.13) $2302.00\pm 0.47$.08) $2385.63\pm 1.76$.04) $2997.61\pm 8.24$.53) Inc-1 $5.80\pm 0.00$.27) $32.71\pm 0.08$.43) $130.06\pm 0.22$.37) $106.98\pm 0.27$.98) $649.05\pm 1.14$.28) $2580.45\pm 0.47$.63) $2471.89\pm 0.47$.46) $3716.82\pm 11.92$.83) Inc-2 $6.44\pm 0.00$.54) $32.98\pm 0.08$.25) $128.01\pm 0.22$.39) $99.39\pm 0.33$.07) $661.57\pm 1.33$.52) $2489.63\pm 4.44$.47) $2240.50\pm 0.43$.84) $2568.75\pm 0.76$.80) Inc-3 $6.85\pm 0.00$.98) $51.18\pm 0.11$.23) $174.20\pm 0.57$.02) $148.67\pm 0.37$.09) $778.78\pm 1.17$.23) $3038.25\pm 0.65$.40) $2930.43\pm 4.30$.59) $3602.56\pm 5.55$.32) Mix-1 $6.85\pm 0.00$.72) $51.94\pm 0.09$.59) $209.07\pm 0.28$.40) $150.07\pm 0.36$.64) $911.69\pm 1.65$.22) $3408.18\pm 5.15$.16) $3379.79\pm 1.11$.99) $4321.64\pm 10.94$.92) Mix-2 $5.16\pm 0.00$.36) $35.97\pm 0.04$.38) $168.30\pm 0.38$.75) $142.36\pm 0.34$.54) $716.46\pm 0.53$.19) $3324.16\pm 11.13$.48) $2794.73\pm 3.45$.73) $3876.34\pm 11.41$.15)
Table 9: Total training runtime (seconds) of $\mathcal{N}_{S}$ on cell domain.
Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K HOPSE-M Adj-1 $8.61\pm 0.02$.06) $38.36\pm 0.12$.45) $153.99\pm 0.20$.76) $155.94\pm 0.40$.02) $856.03\pm 2.24$.92) Adj-2 $5.51\pm 0.00$.81) $32.18\pm 0.10$.10) $169.57\pm 0.35$.97) $153.25\pm 0.21$.41) $1013.70\pm 1.20$.74) Adj-3 $5.50\pm 0.00$.41) $33.48\pm 0.06$.54) $175.46\pm 0.22$.23) $153.43\pm 0.17$.47) $922.75\pm 0.89$.20) Inc-1 $6.98\pm 0.01$.24) $41.18\pm 0.15$.30) $144.62\pm 0.29$.66) $192.74\pm 0.45$.31) $952.75\pm 1.66$.32) Inc-2 $6.81\pm 0.00$.88) $36.17\pm 0.09$.11) $154.09\pm 0.55$.24) $135.15\pm 0.29$.51) $827.17\pm 0.77$.46) Inc-3 $7.15\pm 0.01$.07) $44.15\pm 0.16$.85) $234.99\pm 0.25$.73) $157.33\pm 0.30$.17) $934.61\pm 0.49$.49) Mix-1 $8.88\pm 0.01$.71) $30.30\pm 0.02$.29) $163.48\pm 0.27$.28) $222.96\pm 0.59$.41) $1421.85\pm 1.27$.42) Mix-2 $5.54\pm 0.00$.54) $36.30\pm 0.08$.16) $129.82\pm 0.21$.18) $161.53\pm 0.50$.71) $1046.86\pm 0.47$.16) HOPSE-G Adj-1 $5.09\pm 0.00$.93) $26.71\pm 0.04$.79) $88.31\pm 0.11$.27) $95.95\pm 0.28$.16) $683.86\pm 1.25$.70) Adj-2 $4.98\pm 0.00$.47) $31.85\pm 0.05$.43) $96.33\pm 0.19$.91) $93.98\pm 0.15$.00) $887.83\pm 1.60$.74) Adj-3 $6.64\pm 0.00$.89) $31.13\pm 0.05$.56) $100.65\pm 0.14$.14) $113.39\pm 0.18$.75) $595.39\pm 1.07$.46) Inc-1 $5.75\pm 0.01$.45) $34.11\pm 0.07$.89) $110.19\pm 0.19$.33) $105.67\pm 0.27$.00) $767.42\pm 0.85$.39) Inc-2 $7.65\pm 0.02$.32) $24.65\pm 0.06$.74) $105.02\pm 0.28$.61) $126.14\pm 0.37$.05) $721.57\pm 0.60$.10) Inc-3 $7.87\pm 0.01$.47) $44.42\pm 0.09$.81) $181.18\pm 0.96$.45) $100.22\pm 0.11$.50) $896.86\pm 1.62$.01) Mix-1 $7.61\pm 0.00$.68) $46.14\pm 0.10$.02) $239.59\pm 0.38$.24) $168.35\pm 0.50$.01) $671.40\pm 9.48$.29) Mix-2 $5.27\pm 0.00$.57) $36.29\pm 0.13$.05) $152.73\pm 0.08$.26) $104.12\pm 0.37$.76) $1229.56\pm 5.27$.27)
I.2 Runtime per epoch
Tables˜10 and 11 validate that smaller neighborhoods, and thus less parameters, result in better per epoch runtimes. HOPSE-G has lower per epoch runtime due to having to learn less instantiations of $f_{r,k}$ since $k=1$. Simpler configurations such as Adj-1 consistently yield the lowest runtimes, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets. Note that as dataset size increases, the epoch time scales linearly, thus while HOPSE-G remains the fastest, it is only so by a increasingly smaller fraction.
Table 10: Per epoch runtime (seconds) of $\mathcal{N}_{S}$ on simplicial domain.